OpenStudy (anonymous):

Solve the equation: (x + 6)^3/5 = 8 I know the answer, but not how to find it.

OpenStudy (anonymous):

Just isolate x

OpenStudy (anonymous):

to solve the equation

OpenStudy (anonymous):

not sure how to do that with this equation

OpenStudy (anonymous):

\[(x+6)^\frac{3}{5}=8\\ x+6=8^{\frac{5}{3}}\]

OpenStudy (anonymous):

still not sure what do do from there because the answer is 26 and im not sure how you would get to that

OpenStudy (welshfella):